OpenStudy (anonymous):

4x²+9y²=36; solve for y, then differentiate to get y' in terms of x How do I even begin ?

mathslover (mathslover):

Well , first of all transpose \(4x^2\) to Right Hand side. Can you do that ?

OpenStudy (dan815):

YES can YOU!

OpenStudy (anonymous):

Yup

OpenStudy (dan815):

ok then will you!

OpenStudy (dan815):

mr BURHAN

mathslover (mathslover):

Please do that burhan ... for us

OpenStudy (dan815):

haha ur name is awesome

OpenStudy (anonymous):

okay haha

mathslover (mathslover):

You will get : \(9x^2 = 36-4x^2\) , right?

OpenStudy (anonymous):

\[\huge 36-9y^3=4x^2\]

mathslover (mathslover):

\(9y^2 = 36-4x^2\) I meant this. @burhan101 from where did \(9y^3\) come from, in the question it is \(9y^2\)

OpenStudy (anonymous):

Typo *

OpenStudy (dan815):

haha i was laughing, one guy turns the y to x the other makes it y^3 lol

mathslover (mathslover):

ok. now , divide both sides by 9 : \(9y^2 = 36-4x^2\) \(\cfrac{9y^2}{9} = \cfrac{36-4x^2}{9}\) \(\cfrac{\cancel{9} y^2}{\cancel{9}} = \cfrac{36-4x^2}{9} \) \(y^2 =\cfrac{36-4x^2}{9} \)

OpenStudy (anonymous):

so we're solving for y

mathslover (mathslover):

Right.

mathslover (mathslover):

I have \(y^2\) , now take square root both sides to get \(y\) in terms of x. Can you do that ?

OpenStudy (anonymous):